Kaycie's company pays $20,000 per year plus a 5% sales commission. Kristen's company pays $32,000 per year plus a 1% sales commission.
A.What system of equation can be used to determine the amount Kaycie and Kristen must sell, s, to receive the same pay,p?
B. How much must Kaycie and Kristen sell to receive the same pay? How much will they get paid?

Answer by rfer(16322)

20000+0.05x=32000+0.01x
0.05x-0.01x32000-20000
0.04x=12000
x=12000/0.04
x=300000 in sales